# LibreLog: Accurate and Efficient Unsupervised Log Parsing Using Open-Source Large Language Models

### v2

@@ Abstract
Log parsing is a critical step that transforms unstructured log data into structured formats, facilitating subsequent log-based analysis. Traditional syntax-based log parsers are efficient and effective, but they often experience decreased accuracy when processing logs that deviate from the predefined rules. Recently, large language models (LLM) based log parsers have shown superior parsing accuracy. However, existing LLM-based parsers face three main challenges: 1) time-consuming and labor-intensive manual labeling for fine-tuning or in-context learning, 2) increased parsing costs due to the vast volume of log data and limited context size of LLMs, and 3) privacy risks from using commercial models like ChatGPT with sensitive log information. To overcome these limitations, this paper introduces OpenLogParser, an unsupervised log parsing approach that leverages open-source LLMs (i.e., Llama3-8B) to enhance privacy and reduce operational costs while achieving state-of-the-art parsing accuracy. OpenLogParser first groups logs with similar static text but varying dynamic variables using a fixed-depth grouping tree. It then parses logs within these groups using three components: i) similarity scoring-based retrieval augmented generation: selects diverse logs within each group based on Jaccard similarity, helping the LLM distinguish between static text and dynamic variables; ii) self-reflection: iteratively query LLMs to refine log templates to improve parsing accuracy; and iii) log template memory: stores parsed templates to reduce LLM queries for improved parsing efficiency. Our evaluation on LogHub-2.0 shows that OpenLogParser achieves 25% higher parsing accuracy and processes logs 2.7 times faster compared to state-of-the-art LLM-based parsers. In short, OpenLogParser addresses privacy and cost concerns of using commercial LLMs while achieving state-of-the-arts parsing efficiency and accuracy.
@@ I Introduction
To address these challenges, we propose an unsupervised log parsing technique, OpenLogParser, which does not need any manual labels. OpenLogParser leverages smaller-size open-source LLMs (e.g., Llama3-8B [4]) to enhance privacy and reduce operational costs while achieving state-of-the-art parsing accuracy and efficiency. Inspired by the effective grouping capabilities of syntax-based unsupervised log parsing methods [14], OpenLogParser first groups logs that share syntactic similarity in the static text, but vary in the dynamic variable, using a fixed-depth grouping tree. Then, OpenLogParser parses logs within individual groups through three key steps: (i) OpenLogParser uses similarity scoring-based retrieval augmented generation (RAG) to select the most diverse logs based on Jaccard similarity within each log group. This step helps LLMs separate dynamic and static text by highlighting variability in dynamic variables among logs in the same group. (ii) OpenLogParser uses self-reflection [38] to improve LLM responses, thereby improving parsing results. iii) OpenLogParser uses log template memory to store parsed log templates. This approach allows logs to be parsed by first matching them with stored templates, minimizing the number of LLM queries and significantly enhancing parsing efficiency.
@@
  • We introduce, OpenLogParser, an unsupervised log parsing technique that effectively addresses the limitations of existing LLM-based and syntax-based parsers.
  • OpenLogParser employs open-source LLMs, specifically Llama3-8B, to enhance data privacy and reduce operational costs associated with commercial models.
  • Through extensive evaluations on over 50 million logs from LogHub2.0 [19], OpenLogParser demonstrated a 25% or higher parsing accuracy compared to state-of-the-art LLM-based log parsers (i.e., LILAC [20] and LLMParser [31]). Moreover, it is 2.75 to 40 times faster, showcasing its superior efficiency and effectiveness.
  • OpenLogParser’s self-reflection mechanism helps improve parsing accuracy by over 7%, showcasing the effectiveness of our prompting technique.
@@
Paper Organization. Section II discusses background and related work. Section III provides the design details of OpenLogParser. Section IV outlines evaluation setup. Section V presents evaluation results. Section VI discusses threats to validity. Section VII concludes the paper.
@@ II-B Related Work
In this paper, we propose OpenLogParser that addresses the two above-mentioned limitations. We deployed a relatively small open-source LLM (i.e., Llama3-8B [4]) on log parsing to avoid privacy issues and monetary costs. Additionally, OpenLogParser enhances LLM-based log parsing by capitalizing on the commonalities and variabilities within logs to provide a demonstration-free prompt for the LLM.
@@ III Approach
In this section, we introduce OpenLogParser, an efficient unsupervised log parser, leveraging memory capabilities and advanced prompting techniques to maximize efficiency and parsing accuracy. OpenLogParser leverages a smaller-size open-source LLM to enhance privacy and reduce operation costs. Figure 2 illustrates the overall architecture of OpenLogParser, which primarily comprises of three components: (i) log grouping, which groups logs that share a commonality in their text. Such log groups can then be used as input to LLM to uncover dynamic variables. (ii) An unsupervised LLM-based log parser that uses retrieval-augmented generation (RAG), followed by an iterative self-reflection mechanism to accurately parse the grouped logs into log templates. (iii) An efficient log template memory, which memorizes the parsed log templates for future query. The core idea is to enhance efficiency by storing parsed log templates in memory, thereby avoiding the need for repeated LLM queries.
Figure 2: An overview of OpenLogParser.
@@ III-A Log Grouping Based on Commonality
OpenLogParser achieves unsupervised and zero-shot log parsing by first applying an effective grouping strategy. This strategy aims to group logs that share commonality in their static text, yet are different in their dynamic variables. Such log groups can then be used as input to LLMs to generate log templates by prompting LLMs to identify the dynamic variables among logs in the same group. To group the logs, we adapt the efficient unsupervised methodology proposed by Drain [14], which applies a fixed-depth parsing tree and parsing rules (i.e., $K$ prefix tokens) to identify log groups. The fixed depth in our grouping tree provides a structured and predictable framework that enhances efficiency. By limiting the depth, we reduce the complexity of the tree traversal, which speeds up the grouping process.
@@ III-B LLM-based Unsupervised Log Parsing
Our prompts to LLMs contain representative logs (based on variability) retrieved from each log group (from Section III-A) to guide LLMs in separating dynamic variables and static text. Figure 3 illustrates the prompt template that OpenLogParser uses. Below, we discuss the composition of our prompt in detail.
@@
Figure 3: Example of the prompt template used for OpenLogParser. The green block illustrates the task instruction provided to the LLM. The blue block highlights the input and output examples used to standardize the log response format. The yellow block depicts the retrieval-augmented selection process that enhances log parsing accuracy by incorporating representative variability.
@@ III-C Template Memory for Efficient Log Parsing
Repeatedly using LLM to parse logs with identical groupings and templates significantly increases the frequency of LLM queries, thereby reducing the efficiency of the log parsing process. To address this issue, we introduce log template memory in OpenLogParser, which stores the parsed log templates for future parsing, avoiding redundant LLM queries.
@@ IV Experiment Setup
In this section, we discuss our experiment setup to answer our research questions and OpenLogParser’s implementation details.
@@
Environment and Implementation. Our experiments were conducted on an Ubuntu server with an NVIDIA Tesla A100 GPU, AMD EPYC 7763 64-core CPU, and 256GB RAM using Python 3.9. We execute the baselines using their default parameters under the same environment to compare the efficiency. We use Llama3 8B [4] for OpenLogParser’s underlying LLM because it is a relatively small yet powerful model, balancing performance and efficiency effectively. We set the temperature value to 0 to improve the stability of the model output. Note that it is easy to switch to other LLMs. In RQ4, we evaluate OpenLogParser by replacing Llama3 with other open-source LLMs.
@@ V Evaluation
In this section, we evaluate OpenLogParser by answering four research questions (RQs).
RQ1: What is the effectiveness of OpenLogParser?
Motivation. Accuracy is the most critical factor for evaluating the effectiveness of log parsers. High accuracy in log parsing aids downstream log analysis tasks [22, 37]. In this RQ, we study the effectiveness of OpenLogParser.
TABLE I: A comparison of the grouping accuracy (GA) and parsing accuracy (PA) for the state-of-the-art parsers and OpenLogParser.
     AEL      Drain      LILAC      LLMParserT5Base             OpenLogParser
@@
Approach. We compare OpenLogParser with other state-of-the-art log parsers, including AEL, Drain, LILAC, and LLMParserT5Base. AEL [18] and Drain [14] are two leading traditional syntax-based approaches that are efficient and perform better than most other syntax-based parsers [48, 19]. LILAC [20] and LLMParserT5Base [31] are recently proposed LLM-based parsers with high parsing accuracy. Since LILAC uses ChatGPT as the underlying LLM, for a fair comparison, we replace ChatGPT with the same open-source LLM (Llama3-8B [4]) that OpenLogParser uses. We use T5-base [10] (240M parameters) as the LLM for LLMParser by following the prior work. Note that both LILAC and LLMParser require manually derived log templates as a few shot demonstrations. We follow the steps described in the papers to obtain these demonstrations. We evaluate the parsers using the LogHub-2.0 dataset and report both Grouping Accuracy (GA) and Parsing Accuracy (PA).
@@
Results. Table I shows the GA and PA for each log parser across different systems. OpenLogParser achieved the highest GA and PA values for most systems, indicating superior performance in both grouping and parsing logs. Across all systems, OpenLogParser achieved an average GA of 0.8720 and an average PA of 0.8538, outperforming all other parsers.
@@
OpenLogParser shows superior GA and PA compared to the semi-supervised LLM-based parser – LILAC. Compared to OpenLogParser, LILAC demonstrated lower performance with a GA of 0.8289 and PA of 0.6783. LILAC uses manually labeled logs as demonstrations for in-context learning to enhance parsing accuracy. However, when utilizing less powerful open-source LLMs with smaller parameter sizes (i.e., as opposed to ChatGPT), LILAC’s performance declines significantly due to the limited ability of these models to capture complex log patterns with only a few demonstrations. Consequently, this can lead to inaccurate parsing of variables within the logs (a PA of 0.6783, while OpenLogParser’s PA is 0.8538). Unlike LILAC and LLMParserT5Base, OpenLogParser is an unsupervised log parser, eliminating the need for labeled logs to enhance the LLM’s log parsing capabilities. The performance of OpenLogParser is not dependent on the number of labeled logs, thus avoiding the limitations faced by semi-supervised approaches that require labeled logs for fine-tuning or in-context learning.
@@
Among all three LLM-based log parsers, LLMParserT5Base shows the lowest GA, and the reason may be the limited number of fine-tuning samples that makes it hard to generalize to large-scale datasets. Among the three LLM-based log parsers (LLMParserT5Base, LILAC, and OpenLogParser), LLMParserT5Base exhibited the lowest GA of 0.6345 and the second-highest PA of 0.7514. When parsing large-scale datasets, logs may exhibit many different variations, even if they share the same log template. Given that LLMParserT5Base is fine-tuned using a small, labeled sample set from the target system, the limited number of log samples likely contributes to its inability to robustly identify logs with the same template across all instances and, thus, lower GA. This limitation becomes particularly evident in systems with more logs, such as BGL and Spark, where LLMParserT5Base struggles to achieve high GA (0.1439 and 0.2589, respectively). Nevertheless, it is still able to identify all dynamic variables in a log with the second-highest PA among all five parsers, which shows the potentials of LLM-based parsers.
@@
Syntax-based log parsers generally have significantly lower PAs compared to LLM-based parsers, showing challenges in accurately identifying variables. While AEL and Drain, as syntax-based parsers, show results similar to each other, they both exhibit lower GA compared to OpenLogParser (1.84% and 3.3% lower, respectively) and significantly lower PA (48.25% and 45.17% lower, respectively). This performance disparity is likely linked to their heuristic-based nature, which relies on predefined rules to identify log features. While these rules can effectively classify logs with similar features, achieving reasonable GAs, their generic nature often fails to accurately recognize variables within different log templates, leading to poor PAs. In contrast, OpenLogParser leverages pre-grouping and uses memory mechanisms to achieve high GA, and its LLM-based parsing process accurately identifies variables within grouped logs, resulting in superior PA.
@@
OpenLogParser achieves superior GA and PA compared to state-of-the-art parsers. Despite not relying on labeled logs, OpenLogParser outperforms other LLM-based parsers that are semi-supervised. Additionally, OpenLogParser significantly enhances PA compared to syntax-based approaches.
RQ2: What is the efficiency of OpenLogParser?
@@
TABLE II: Number of logs and parsing time, in seconds, for the state-of-the-art (first four columns) and OpenLogParser.
AEL Drain LILAC LLMParserT5Base     OpenLogParser
@@
Approach. We measure the total parsing time required by OpenLogParser and its individual components (i.e., LLM queries, grouping, and memory search), and the four baseline parsers to process logs from the LogHub-2.0 dataset.
@@
Results. OpenLogParser is 2.7 and 40 times faster than Lilac and LLMParserT5Base, respectively. Table II shows the parsing time for each log parser across different systems. OpenLogParser spends a total of 5.94 hours to parse logs from all 14 systems (50 million logs), which is significantly faster than other LLM-based parsers: LILAC (16 hours) and LLMParserT5Base (258 hours). The parsing time for OpenLogParser is mainly occupied by the LLM query time, which accounts for 72.05% of the total processing time, followed by the grouping time, which constitutes 16.67% of the overall duration. LLMParserT5Base is the slowest among all LLM-based parsers because it processes each log individually, and the vast quantity of logs linearly increases the number of model queries required. Even with a relatively lightweight model like T5-base, which has only 240 million parameters, querying to parse the logs individually is still slow and impractical for real-world applications. LILAC, with its cache design, eliminates the need to parse each log individually through an LLM, significantly speeding up the process compared to LLMParserT5Base. However, LILAC still requires frequent model queries to update the templates in the cache, which limits its efficiency. In contrast, OpenLogParser optimizes parsing times through its grouping and memory features, resulting in superior efficiency.
@@
AEL exhibits significant efficiency issues when parsing logs beyond certain sizes, while Drain maintains high efficiency across all datasets. AEL can parse datasets with fewer than 100K logs within seconds but requires several hours or even days for datasets with over one million logs (e.g., we stopped AEL after running for 10 days when parsing the 16 million logs from Spark). This inefficiency is due to AEL’s reliance on extensive comparisons between logs and identified templates, where the parsing time grows exponentially with respect to the number of logs and log templates. In contrast, Drain, which uses a fixed-depth parsing tree, is the most efficient parser. OpenLogParser uses a grouping method similar to Drain’s, with a total grouping time amounting to 0.99 hours, which is less than Drain’s total parsing time of 1.6 hours. This highlights the efficiency of OpenLogParser’s grouping process. While there is a slight slowdown due to the additional processing involved (5.94 hours compared to Drain’s 1.6 hours), OpenLogParser shows superior parsing effectiveness compared to Drain and is the second fastest log parser among the evaluates parsers.
@@
OpenLogParser enhances its efficiency by utilizing grouping and memory components, which reduces the number of LLM queries. OpenLogParser demonstrates the highest efficiency across LLM-based parsers.
RQ3: How does different settings impact the result of OpenLogParser?
@@
Motivation. OpenLogParser implements multiple components to achieve effective and efficient log parsing. In this RQ, we explore how various settings and configurations affect the performance of OpenLogParser.
@@
Approach. There are three general components in OpenLogParser that can be adjusted or replaced: log selection from each group for prompting, the number of selected logs, and the inclusion or exclusion of self-reflection processes. To select diverse logs from the log group, we use Jaccard similarity to measure the similarity between every log pair. In this RQ, we also try random sampling and cosine similarity. Furthermore, we evaluate how changing the number of selected logs from 1 to 10 impacts the effectiveness. Finally, we compare the effect of removing the self-reflection component on the efficiency and effectiveness of OpenLogParser.
@@
Results. Selecting representative logs based on Jaccard similarity outperforms using cosine similarity and random sampling. Table III shows the total time, GA, and PA of OpenLogParser compared to replacing the log selection process with cosine similarity and random sampling. When employing cosine similarity to select representative logs, both GA and PA experienced declines of 2.5% and 4.8%, respectively, compared to using Jaccard similarity. This indicates that although cosine similarity is shown to be an effective similarity metric for text data [39], it does not necessarily select logs that are representative enough for LLM to generalize log templates. However, we notice a slight reduction in execution time (3.3%) when using cosine similarity. Similarly, using random sampling further reduces the processing time (by 8.2%), but due to the lack of diversity in the sampled logs, both GA and PA are even lower, at 0.849 and 0.806, respectively.
@@
TABLE III: OpenLogParser performance under different settings. The numbers in the parenthesis indicate the percentage difference compared to the full version of OpenLogParser.
@@
Although the self-reflection mechanism requires additional processing time, it significantly enhances the parsing results of OpenLogParser. Table III compares full version OpenLogParser and OpenLogParser without self-reflection in the total execution time, GA, and PA. Excluding the self-reflection component from OpenLogParser results in a 44.6% reduction in parsing time (from around six to three hours). However, removing self-reflection greatly decreases both GA and PA by 7.1% and 9.5%, respectively. This shows that self-reflection significantly enhances the parsing effectiveness of OpenLogParser, although at the expense of increased overhead due to additional LLM queries. Therefore, in practical applications, the inclusion of the self-reflection component in OpenLogParser can be determined based on the specific needs of effectiveness or efficiency.

### v3

@@ Abstract
Log parsing is a critical step that transforms unstructured log data into structured formats, facilitating subsequent log-based analysis. Traditional syntax-based log parsers are efficient and effective, but they often experience decreased accuracy when processing logs that deviate from the predefined rules. Recently, large language models (LLM) based log parsers have shown superior parsing accuracy. However, existing LLM-based parsers face three main challenges: 1) time-consuming and labor-intensive manual labeling for fine-tuning or in-context learning, 2) increased parsing costs due to the vast volume of log data and limited context size of LLMs, and 3) privacy risks from using commercial models like ChatGPT with sensitive log information. To overcome these limitations, this paper introduces LibreLog, an unsupervised log parsing approach that leverages open-source LLMs (i.e., Llama3-8B) to enhance privacy and reduce operational costs while achieving state-of-the-art parsing accuracy. LibreLog first groups logs with similar static text but varying dynamic variables using a fixed-depth grouping tree. It then parses logs within these groups using three components: i) similarity scoring-based retrieval augmented generation: selects diverse logs within each group based on Jaccard similarity, helping the LLM distinguish between static text and dynamic variables; ii) self-reflection: iteratively query LLMs to refine log templates to improve parsing accuracy; and iii) log template memory: stores parsed templates to reduce LLM queries for improved parsing efficiency. Our evaluation on LogHub-2.0 shows that LibreLog achieves 25% higher parsing accuracy and processes logs 2.7 times faster compared to state-of-the-art LLM-based parsers. In short, LibreLog addresses privacy and cost concerns of using commercial LLMs while achieving state-of-the-arts parsing efficiency and accuracy.
@@ I Introduction
To address these challenges, we propose an unsupervised log parsing technique, LibreLog, which does not need any manual labels. LibreLog leverages smaller-size open-source LLMs (e.g., Llama3-8B [4]) to enhance privacy and reduce operational costs while achieving state-of-the-art parsing accuracy and efficiency. Inspired by the effective grouping capabilities of syntax-based unsupervised log parsing methods [14], LibreLog first groups logs that share syntactic similarity in the static text, but vary in the dynamic variable, using a fixed-depth grouping tree. Then, LibreLog parses logs within individual groups through three key steps: (i) LibreLog uses similarity scoring-based retrieval augmented generation (RAG) to select the most diverse logs based on Jaccard similarity within each log group. This step helps LLMs separate dynamic and static text by highlighting variability in dynamic variables among logs in the same group. (ii) LibreLog uses self-reflection [38] to improve LLM responses, thereby improving parsing results. iii) LibreLog uses log template memory to store parsed log templates. This approach allows logs to be parsed by first matching them with stored templates, minimizing the number of LLM queries and significantly enhancing parsing efficiency.
@@
  • We introduce, LibreLog, an unsupervised log parsing technique that effectively addresses the limitations of existing LLM-based and syntax-based parsers.
  • LibreLog employs open-source LLMs, specifically Llama3-8B, to enhance data privacy and reduce operational costs associated with commercial models.
  • Through extensive evaluations on over 50 million logs from LogHub2.0 [19], LibreLog demonstrated a 25% or higher parsing accuracy compared to state-of-the-art LLM-based log parsers (i.e., LILAC [20] and LLMParser [31]). Moreover, it is 2.75 to 40 times faster, showcasing its superior efficiency and effectiveness.
  • LibreLog’s self-reflection mechanism helps improve parsing accuracy by over 7%, showcasing the effectiveness of our prompting technique.
@@
Paper Organization. Section II discusses background and related work. Section III provides the design details of LibreLog. Section IV outlines evaluation setup. Section V presents evaluation results. Section VI discusses threats to validity. Section VII concludes the paper.
@@ II-B Related Work
In this paper, we propose LibreLog that addresses the two above-mentioned limitations. We deployed a relatively small open-source LLM (i.e., Llama3-8B [4]) on log parsing to avoid privacy issues and monetary costs. Additionally, LibreLog enhances LLM-based log parsing by capitalizing on the commonalities and variabilities within logs to provide a demonstration-free prompt for the LLM.
@@ III Approach
In this section, we introduce LibreLog, an efficient unsupervised log parser, leveraging memory capabilities and advanced prompting techniques to maximize efficiency and parsing accuracy. LibreLog leverages a smaller-size open-source LLM to enhance privacy and reduce operation costs. Figure 2 illustrates the overall architecture of LibreLog, which primarily comprises of three components: (i) log grouping, which groups logs that share a commonality in their text. Such log groups can then be used as input to LLM to uncover dynamic variables. (ii) An unsupervised LLM-based log parser that uses retrieval-augmented generation (RAG), followed by an iterative self-reflection mechanism to accurately parse the grouped logs into log templates. (iii) An efficient log template memory, which memorizes the parsed log templates for future query. The core idea is to enhance efficiency by storing parsed log templates in memory, thereby avoiding the need for repeated LLM queries.
Figure 2: An overview of LibreLog.
@@ III-A Log Grouping Based on Commonality
LibreLog achieves unsupervised and zero-shot log parsing by first applying an effective grouping strategy. This strategy aims to group logs that share commonality in their static text, yet are different in their dynamic variables. Such log groups can then be used as input to LLMs to generate log templates by prompting LLMs to identify the dynamic variables among logs in the same group. To group the logs, we adapt the efficient unsupervised methodology proposed by Drain [14], which applies a fixed-depth parsing tree and parsing rules (i.e., $K$ prefix tokens) to identify log groups. The fixed depth in our grouping tree provides a structured and predictable framework that enhances efficiency. By limiting the depth, we reduce the complexity of the tree traversal, which speeds up the grouping process.
@@ III-B LLM-based Unsupervised Log Parsing
Our prompts to LLMs contain representative logs (based on variability) retrieved from each log group (from Section III-A) to guide LLMs in separating dynamic variables and static text. Figure 3 illustrates the prompt template that LibreLog uses. Below, we discuss the composition of our prompt in detail.
@@
Figure 3: Example of the prompt template used for LibreLog. The green block illustrates the task instruction provided to the LLM. The blue block highlights the input and output examples used to standardize the log response format. The yellow block depicts the retrieval-augmented selection process that enhances log parsing accuracy by incorporating representative variability.
@@ III-C Template Memory for Efficient Log Parsing
Repeatedly using LLM to parse logs with identical groupings and templates significantly increases the frequency of LLM queries, thereby reducing the efficiency of the log parsing process. To address this issue, we introduce log template memory in LibreLog, which stores the parsed log templates for future parsing, avoiding redundant LLM queries.
@@ IV Experiment Setup
In this section, we discuss our experiment setup to answer our research questions and LibreLog’s implementation details.
@@
Environment and Implementation. Our experiments were conducted on an Ubuntu server with an NVIDIA Tesla A100 GPU, AMD EPYC 7763 64-core CPU, and 256GB RAM using Python 3.9. We execute the baselines using their default parameters under the same environment to compare the efficiency. We use Llama3 8B [4] for LibreLog’s underlying LLM because it is a relatively small yet powerful model, balancing performance and efficiency effectively. We set the temperature value to 0 to improve the stability of the model output. Note that it is easy to switch to other LLMs. In RQ4, we evaluate LibreLog by replacing Llama3 with other open-source LLMs.
@@ V Evaluation
In this section, we evaluate LibreLog by answering four research questions (RQs).
RQ1: What is the effectiveness of LibreLog?
Motivation. Accuracy is the most critical factor for evaluating the effectiveness of log parsers. High accuracy in log parsing aids downstream log analysis tasks [22, 37]. In this RQ, we study the effectiveness of LibreLog.
TABLE I: A comparison of the grouping accuracy (GA) and parsing accuracy (PA) for the state-of-the-art parsers and LibreLog.
     AEL      Drain      LILAC      LLMParserT5Base             LibreLog
@@
Approach. We compare LibreLog with other state-of-the-art log parsers, including AEL, Drain, LILAC, and LLMParserT5Base. AEL [18] and Drain [14] are two leading traditional syntax-based approaches that are efficient and perform better than most other syntax-based parsers [48, 19]. LILAC [20] and LLMParserT5Base [31] are recently proposed LLM-based parsers with high parsing accuracy. Since LILAC uses ChatGPT as the underlying LLM, for a fair comparison, we replace ChatGPT with the same open-source LLM (Llama3-8B [4]) that LibreLog uses. We use T5-base [10] (240M parameters) as the LLM for LLMParser by following the prior work. Note that both LILAC and LLMParser require manually derived log templates as a few shot demonstrations. We follow the steps described in the papers to obtain these demonstrations. We evaluate the parsers using the LogHub-2.0 dataset and report both Grouping Accuracy (GA) and Parsing Accuracy (PA).
@@
Results. Table I shows the GA and PA for each log parser across different systems. LibreLog achieved the highest GA and PA values for most systems, indicating superior performance in both grouping and parsing logs. Across all systems, LibreLog achieved an average GA of 0.8720 and an average PA of 0.8538, outperforming all other parsers.
@@
LibreLog shows superior GA and PA compared to the semi-supervised LLM-based parser – LILAC. Compared to LibreLog, LILAC demonstrated lower performance with a GA of 0.8289 and PA of 0.6783. LILAC uses manually labeled logs as demonstrations for in-context learning to enhance parsing accuracy. However, when utilizing less powerful open-source LLMs with smaller parameter sizes (i.e., as opposed to ChatGPT), LILAC’s performance declines significantly due to the limited ability of these models to capture complex log patterns with only a few demonstrations. Consequently, this can lead to inaccurate parsing of variables within the logs (a PA of 0.6783, while LibreLog’s PA is 0.8538). Unlike LILAC and LLMParserT5Base, LibreLog is an unsupervised log parser, eliminating the need for labeled logs to enhance the LLM’s log parsing capabilities. The performance of LibreLog is not dependent on the number of labeled logs, thus avoiding the limitations faced by semi-supervised approaches that require labeled logs for fine-tuning or in-context learning.
@@
Among all three LLM-based log parsers, LLMParserT5Base shows the lowest GA, and the reason may be the limited number of fine-tuning samples that makes it hard to generalize to large-scale datasets. Among the three LLM-based log parsers (LLMParserT5Base, LILAC, and LibreLog), LLMParserT5Base exhibited the lowest GA of 0.6345 and the second-highest PA of 0.7514. When parsing large-scale datasets, logs may exhibit many different variations, even if they share the same log template. Given that LLMParserT5Base is fine-tuned using a small, labeled sample set from the target system, the limited number of log samples likely contributes to its inability to robustly identify logs with the same template across all instances and, thus, lower GA. This limitation becomes particularly evident in systems with more logs, such as BGL and Spark, where LLMParserT5Base struggles to achieve high GA (0.1439 and 0.2589, respectively). Nevertheless, it is still able to identify all dynamic variables in a log with the second-highest PA among all five parsers, which shows the potentials of LLM-based parsers.
@@
Syntax-based log parsers generally have significantly lower PAs compared to LLM-based parsers, showing challenges in accurately identifying variables. While AEL and Drain, as syntax-based parsers, show results similar to each other, they both exhibit lower GA compared to LibreLog (1.84% and 3.3% lower, respectively) and significantly lower PA (48.25% and 45.17% lower, respectively). This performance disparity is likely linked to their heuristic-based nature, which relies on predefined rules to identify log features. While these rules can effectively classify logs with similar features, achieving reasonable GAs, their generic nature often fails to accurately recognize variables within different log templates, leading to poor PAs. In contrast, LibreLog leverages pre-grouping and uses memory mechanisms to achieve high GA, and its LLM-based parsing process accurately identifies variables within grouped logs, resulting in superior PA.
@@
LibreLog achieves superior GA and PA compared to state-of-the-art parsers. Despite not relying on labeled logs, LibreLog outperforms other LLM-based parsers that are semi-supervised. Additionally, LibreLog significantly enhances PA compared to syntax-based approaches.
RQ2: What is the efficiency of LibreLog?
@@
TABLE II: Number of logs and parsing time, in seconds, for the state-of-the-art (first four columns) and LibreLog.
AEL Drain LILAC LLMParserT5Base     LibreLog
@@
Approach. We measure the total parsing time required by LibreLog and its individual components (i.e., LLM queries, grouping, and memory search), and the four baseline parsers to process logs from the LogHub-2.0 dataset.
@@
Results. LibreLog is 2.7 and 40 times faster than Lilac and LLMParserT5Base, respectively. Table II shows the parsing time for each log parser across different systems. LibreLog spends a total of 5.94 hours to parse logs from all 14 systems (50 million logs), which is significantly faster than other LLM-based parsers: LILAC (16 hours) and LLMParserT5Base (258 hours). The parsing time for LibreLog is mainly occupied by the LLM query time, which accounts for 72.05% of the total processing time, followed by the grouping time, which constitutes 16.67% of the overall duration. LLMParserT5Base is the slowest among all LLM-based parsers because it processes each log individually, and the vast quantity of logs linearly increases the number of model queries required. Even with a relatively lightweight model like T5-base, which has only 240 million parameters, querying to parse the logs individually is still slow and impractical for real-world applications. LILAC, with its cache design, eliminates the need to parse each log individually through an LLM, significantly speeding up the process compared to LLMParserT5Base. However, LILAC still requires frequent model queries to update the templates in the cache, which limits its efficiency. In contrast, LibreLog optimizes parsing times through its grouping and memory features, resulting in superior efficiency.
@@
AEL exhibits significant efficiency issues when parsing logs beyond certain sizes, while Drain maintains high efficiency across all datasets. AEL can parse datasets with fewer than 100K logs within seconds but requires several hours or even days for datasets with over one million logs (e.g., we stopped AEL after running for 10 days when parsing the 16 million logs from Spark). This inefficiency is due to AEL’s reliance on extensive comparisons between logs and identified templates, where the parsing time grows exponentially with respect to the number of logs and log templates. In contrast, Drain, which uses a fixed-depth parsing tree, is the most efficient parser. LibreLog uses a grouping method similar to Drain’s, with a total grouping time amounting to 0.99 hours, which is less than Drain’s total parsing time of 1.6 hours. This highlights the efficiency of LibreLog’s grouping process. While there is a slight slowdown due to the additional processing involved (5.94 hours compared to Drain’s 1.6 hours), LibreLog shows superior parsing effectiveness compared to Drain and is the second fastest log parser among the evaluates parsers.
@@
LibreLog enhances its efficiency by utilizing grouping and memory components, which reduces the number of LLM queries. LibreLog demonstrates the highest efficiency across LLM-based parsers.
RQ3: How does different settings impact the result of LibreLog?
@@
Motivation. LibreLog implements multiple components to achieve effective and efficient log parsing. In this RQ, we explore how various settings and configurations affect the performance of LibreLog.
@@
Approach. There are three general components in LibreLog that can be adjusted or replaced: log selection from each group for prompting, the number of selected logs, and the inclusion or exclusion of self-reflection processes. To select diverse logs from the log group, we use Jaccard similarity to measure the similarity between every log pair. In this RQ, we also try random sampling and cosine similarity. Furthermore, we evaluate how changing the number of selected logs from 1 to 10 impacts the effectiveness. Finally, we compare the effect of removing the self-reflection component on the efficiency and effectiveness of LibreLog.
@@
Results. Selecting representative logs based on Jaccard similarity outperforms using cosine similarity and random sampling. Table III shows the total time, GA, and PA of LibreLog compared to replacing the log selection process with cosine similarity and random sampling. When employing cosine similarity to select representative logs, both GA and PA experienced declines of 2.5% and 4.8%, respectively, compared to using Jaccard similarity. This indicates that although cosine similarity is shown to be an effective similarity metric for text data [39], it does not necessarily select logs that are representative enough for LLM to generalize log templates. However, we notice a slight reduction in execution time (3.3%) when using cosine similarity. Similarly, using random sampling further reduces the processing time (by 8.2%), but due to the lack of diversity in the sampled logs, both GA and PA are even lower, at 0.849 and 0.806, respectively.
@@
TABLE III: LibreLog performance under different settings. The numbers in the parenthesis indicate the percentage difference compared to the full version of LibreLog.
@@
Although the self-reflection mechanism requires additional processing time, it significantly enhances the parsing results of LibreLog. Table III compares full version LibreLog and LibreLog without self-reflection in the total execution time, GA, and PA. Excluding the self-reflection component from LibreLog results in a 44.6% reduction in parsing time (from around six to three hours). However, removing self-reflection greatly decreases both GA and PA by 7.1% and 9.5%, respectively. This shows that self-reflection significantly enhances the parsing effectiveness of LibreLog, although at the expense of increased overhead due to additional LLM queries. Therefore, in practical applications, the inclusion of the self-reflection component in LibreLog can be determined based on the specific needs of effectiveness or efficiency.